OpenStudy (anonymous):

Simplify this problem (4d^9)^2/(-2c^5)^6

OpenStudy (anonymous):

\[\frac{( 4d ^{9} )^{2}}{( -2c ^{5^){6}} }\]

OpenStudy (anonymous):

Law of Exponents: \[(A^x)^y=A^{xy}\]\[A^xB^x=(AB)^x\]\[\frac{ (4d^9)^2 }{ (-2c^5)^6 }=\frac{ (4)^2(d^9)^2 }{ (-2)^6(c^5)^6 }=\frac{ 16d^{18} }{ 64c^{30} }=\frac{ d^{18} }{ 4c^{30} }\]

OpenStudy (anonymous):

@C.Hernandez